OpenStudy (anonymous):

how do i find the y- intercept of f(x)= (x^2-6x-7)/(4x^2-36x+56)

OpenStudy (anonymous):

Hey again

OpenStudy (anonymous):

hello

OpenStudy (anonymous):

The prerequisite for an intersection with the \(y\)-axis is \(x\) to be zero, which you can simply plug in :\[ f(x) = \frac{x^2-6x-7}{4x^2-36x+56} \Rightarrow f(0) \frac{0^2-6\cdot 0-7}{4\cdot0^2-36\cdot0+56} \]

OpenStudy (anonymous):

Can you figure out the rest yourself?

OpenStudy (anonymous):

do i leave it in that form?

OpenStudy (anonymous):

no, you calculate the actual values \[ \frac{0-0-7}{0-0+56} = -\frac{7}{56} = -0.125 \]

OpenStudy (anonymous):

i meant did i have to leave it in fraction form?

OpenStudy (anonymous):

i think thats your choice. I would write it down as simple an accurate as possible. As -0.125 is a number with finite digits it's as accurate as it gets

OpenStudy (anonymous):

given: g(x)=x^3-18x^2+85x-156 use division to simplify g into the product of a linear and a quadratic factor

OpenStudy (anonymous):

I gotta go now unless you need to know it very fast you can write me a message and i'll help you as soon as i come back here. Otherwise simply ask it in public

OpenStudy (anonymous):

ok thank you for help

OpenStudy (anonymous):

Again, you're welcome